# Harmonic Mobile Manipulation

### v1

@@ 5.1 Training and Evaluation Details
5.2 Performance of HarmonicMM in Simulation
@@
We trained HarmonicMM and evaluated its performance on four tasks in our Daily Mobile Manipulation Task Suite, each requiring navigation and manipulation skills. Our experiment results demonstrate that HarmonicMM achieves an average success rate of 47% across all tasks (As in Table 1). Specifically, our model attains a high success rate of 80% for the task of Opening Door (Push) in unseen houses, and achieves the average progress rates of 76%, 70%, and 46% for Cleaning Table, Opening Door (Pull), and Opening Fridge, respectively, indicate significant progress in each task. This achievement is especially impressive given the complex, long-horizon nature of these tasks.
@@
The success rates across tasks vary due to how the robot interacts with its environment. Opening Door (Push) has a high success rate as it’s less affected by environmental factors: the robot can use the space created by the opened door to continue the task. Conversely, Opening Door (Pull) and Opening Fridge are more challenging. These tasks reduce the available space as they progress, especially in narrow areas, leading to lower success rates despite significant progress. Cleaning Table involves complex navigation due to obstacles and furniture around the table, requiring intricate movement patterns. This complexity accounts for its lower performance compared to door-related tasks. For Opening Fridge, the task’s difficulty arises from the smaller size of fridge doors and the need for precise end-effector placement. The often confined placement of fridges in rooms adds to the challenge, affecting the robot’s ability to navigate and manipulate effectively.
Figure 6: Cleaning Table in Simulation: Our robot moves towards the table while stretching its arm in simulation.
5.3 Comparison with the previous approaches
@@
Our hypothesis is that jointly learning and executing navigation and manipulation for mobile manipulation tasks achieves better performance than considering them disjointly. To validate this, we compared our method with two common approaches widely used by the community. For a fair comparison, we use the same network architecture as stated in Sec 4 and the same hyperparameters for baselines.
@@
One common practice, which we note as Two-Stage baseline, in embodied AI for mobile manipulation [48, 12, 37] is to decompose the tasks into separate stages of manipulation and navigation. In this work, our Two-Stage baseline uses a roughly similar high-level structure as OVMM[48] that robot navigates to the target object and, after reaching the target, switching to the manipulation. While this approach works for a variety of pick-and-place tasks, it fails for more complex mobile manipulation tasks, such as those presented in our benchmark. Table 1 shows the very low success rate and progress of the Two-Stage baseline.
@@
Another common approach is executing manipulation and navigation alternatively, at every step controller invokes separately obtained manipulation or navigation skills. We adapted ReLMoGen [44] to support the embodiment of the Stretch robot. This baseline outputs the target end-effector pose, target base pose in the current robot frame, and choice over navigation or manipulation at every step. Although this approach achieves much higher results than the Two-Stage baseline, there is still a significant 17.6% (37% relative) drop in success rate compared to HarmonicMM.
5.4 HarmonicMM is Efficient.
One of our main hypotheses was that conducting mobile manipulation in a harmonious approach would not only increase the overall performance of the model but also make the agents more efficient. Consider table cleaning as an example. We lift our hands as we approach the table (as opposed to moving toward the table and then lifting our hands) as the trajectory of our robot in Fig.6 in simulation. Table 2 shows that HarmonicMM makes 32.2% more progress towards completing the task at each step compared to the baselines on Cleaning Table, 113.4% on Opening Door (Push), and 27.6% on Opening Door (Pull).
5.5 HarmonicMM transfers to the real world.
Table 4: Real World Evaluation: We evaluated our method in the real world, demonstrating successful sim2real transfer.
@@
Cleaning Table $62.5\%$ 8
Opening Door (Pull) $60.0\%$ 5
Opening Door (Push) $80.0\%$ 5
We also deployed our learned policy in a real apartment for Opening Door (Pull), Opening Door (Push), and Cleaning Table tasks, as depicted in Fig. 5. The Opening Fridge task is not evaluated since the magnetic seal of the fridge is too strong for our robot to pull. Since we simplified the grasping in simulation using a magnetic gripper, we provided a heuristic function for grasping the doorknob in the real world to bridge this gap. When our controller initiates a grasp action, the heuristic function is invoked to grasp the target object. In the real-world evaluation, our model was tested in a few trials due to the extended duration of each task. Despite this, as shown in Table 4, it achieved notable success. While the small number of trials limits the statistical significance of our results, the controller’s successful performance in these real-world tests is a promising indicator of its potential. This preliminary success demonstrates the effectiveness of our approach in managing the complexities and unpredictabilities of real-world environments.
@@
Our controller successfully pulled the door fully open in 3 out of 5 attempts in three different rooms. Interestingly, we observed that HarmonicMM exhibited two distinct styles of pulling the door to adapt to the layout differences in each room. When there was enough open space around the door, it continued pulling until the door was fully opened, as shown in the 1st row in Fig.5. However, when the door was located next to a wall, the agent first pulled the door to start the opening process and then switched to pushing to fully open it (2nd row, Fig.5). These emergent behaviors showcase the spatial reasoning and scene-understanding abilities of our controller. Similarly, our controller successfully pushed doors fully open in 4 different rooms out of 5 trials.
@@
For the cleaning table task, we placed the table in three different locations across two rooms, with evenly distributed paper pieces on it for the robot to clean. Our robot navigated to the table and continuously moved its end-effector to clean the entire table, as shown in Fig. 5. Out of 8 trials, our controller successfully cleaned 75% of the table’s surface area in five instances.
@@
Figure 7: Improvement over Range: We evaluate the improvement HarmonicMM obtained over ReLMoGen baseline when initialized at positions with different distances towards the target object. The improvement grows as the distance range increases.
5.6 HarmonicMM is more efficient in longer horizon tasks
@@
To better understand the improvements achieved by HarmonicMM over the baselines, we analyzed the success rate of both HarmonicMM and ReLMoGen[44] in the Cleaning Table and Opening Door (Push) tasks across varying initial distances from the target object of interest (aka. table and door), ranging from $1\sim 4$. As illustrated in Fig. 7, we observed that the performance improvement of HarmonicMM over the baseline becomes larger as the initial distance to the target object increases. This trend underscores HarmonicMM’s enhanced efficiency, particularly in tasks with longer horizons, where the coordinated integration of navigation and manipulation plays a pivotal role in successful task completion.
5.7 Proprioception in HarmonicMM
While proprioception is commonly used in robotics, it’s less prevalent in embodied AI tasks. To assess its impact in our pipeline, we compared the HarmonicMM with a vision-only variant (HarmonicMM w/o proprio) on the Opening Door (Push) and Opening Door (Pull) tasks. Table 3 shows a notable performance decline in the Opening Door (Pull) task without proprioception. As detailed in Sec 4.1, pulling a door necessitates pre-grasping the knob, which requires precise end-effector positioning. This finding highlights the efficiency of direct proprioceptive input over relying solely on visual observations for manipulation, especially in tasks where accurate end-effector positioning is crucial. Conversely, in the Opening Door (Push) task, where precise positioning is less critical, we still observed a minor performance drop. This suggests that even in tasks where accurate end-effector placement is not essential, proprioception contributes positively to both performance and efficiency.
6 Discussion
@@
In this work, we developed a simulation for complex mobile manipulation tasks, ranging from opening doors/fridges to cleaning tables in daily scenes, using ProcThor[5]. Our HarmonicMM enables robots to solve these tasks using only RGB visual observation and proprioception of the robot through end-to-end learning. Our pipeline significantly outperforms previous baselines in simulation and has successfully transferred to real world apartments with novel layouts, without any fine-tuning. Code for our benchmark and our method will be released.
Limitation: While our HarmonicMM pipeline allows the Hello Robot Stretch[18] to perform complex mobile manipulation tasks, there are inherent limitations due to the kinematic constraints and the physical capabilities of the robot. These constraints prevent it from completing more dynamic tasks in the real world, such as manipulating heavy objects.
Future Works: In future work, our aim is to enhance the capabilities of HarmonicMM by integrating more complex and dynamic tasks into our task suite. We aim to explore the potential of HarmonicMM in environments with even longer task horizons and more varied challenges. For instance, tasks such as navigating from one room to another while performing sequential tasks like cleaning, object retrieval, or interacting with appliances will be investigated.
@@
Acknowledgements. We extend our gratitude to Rose Hendrix for her invaluable assistance with the hardware setup. We would also like to thank the Thor Team, comprising Winson Han, Eli VanderBilt, and Alvaro Herrasti, for their support in setting up the simulations. Special thanks to Winson Han for his expertise and assistance in visualization of this work.
Supplementary Material
7 Visual Augmentation
To mitigate the simulation-to-reality (sim2real) gap in visual observations, we have implemented a series of augmentations to enhance robustness and generalization, as described in Sec 4. These augmentations are specifically designed to account for variations in visual input that a robot might encounter in real-world scenarios. The applied augmentations include:
  • ColorJitter: Adjusts brightness (0.4), contrast (0.4), saturation (0.2), and hue (0.05) to simulate different lighting conditions and color variations.
  • GaussianBlur: Applies a blur effect with a kernel size of (5, 9) and sigma range of (0.1, 2), mimicking potential blurriness in real-world vision.
  • RandomResizedCrop: Resizes the visual input with a scale range of (0.9, 1) to replicate variations in the field of view and distance.
  • RandomPosterize: Reduces color depth to simulate lower quality images, applied with varying bits (7, 6, 5, 4) and a probability of 0.2 for each setting.
  • RandomAdjustSharpness: Enhances or reduces the sharpness with a factor of 2, applied with a probability of 0.5, to reflect different levels of image clarity.
These augmentations collectively aim to prepare the model for diverse visual scenarios, bridging the gap between simulated and real-world environments.
8 Network Architecture
In this section, we delve into the specifics of our HarmonicMM network architecture as described in Sec 4.3. The architecture processes RGB observations from dual visual cameras, each with a resolution of $384\times 224$. These observations undergo initial processing through a shared visual backbone, specifically the DinoV2 model, which generates latent features of dimensions $768\times 16\times 16$. These features are then adaptively pooled to a reduced dimension of $768\times 7\times 7$. Subsequently, the processed features are passed through separate visual encoders, $E_{Nav}$ for navigation and $E_{Manip}$ for manipulation. This results in two distinct sets of latent visual features, each with a shape of $16\times 7\times 7$, corresponding to each camera view. These visual features are concatenated with proprioceptive observations and input into a single-layer GRU with 512 hidden units. This is followed by a linear layer, also with 512 hidden units, to derive the action or value outputs required for reinforcement learning (RL) training.
The key hyperparameters of our architecture are as follows:
@@ 8 Network Architecture
Visual Backbone dinov2_vits14
Visual Enc $E_{Nav}$ 2 Conv($1\times 1$) layers of 16 channels
Visual Enc $E_{Manip}$ 2 Conv($1\times 1$) layers of16 channels
@@
Policy($\pi$) 1 FC layer with 512 units
Value($v$) 1 FC layer with 512 units
@@
latent feature dim $7\times 7\times 16*2+5=1573$
9 Hyperparameters
Here we provide hyperparameters related to RL training and shared by all methods.
@@ 9 Hyperparameters
Optimizer Adam
10 Reward Function
In this section, we provide detailed formulation of our reward and corresponding hyperparameters for different tasks as described in Sec 4.4.
Our reward function includes the following terms:
$R=w_{\text{nav}}R_{\text{nav}}+w_{\text{manip}}R_{\text{manip}}+w_{\text{efficiency}}R_{\text{efficiency}}$
@@ 10 Reward Function
$R_{\text{nav}}=\delta(w_{\text{nav shaping}}max(d_{\text{cloest}}-d_{\text{current}},0)+\gamma R_{\text{reach target}})$
where $\delta$ is defined as:
$\delta=\begin{cases}1,&\text{if the robot haven't reached the target position}\\ 0,&\text{otherwise}\end{cases}$
$\gamma$ is defined as:
$\gamma=\begin{cases}1,&\text{if the robot reaches the target at the current step}\\ 0,&\text{otherwise}\end{cases}$
$d_{\text{closet}}$ is the closest distance robot reached before, $d_{\text{current}}$ is the current distance between robot and target object. For all tasks, we use $w_{\text{nav shaping}}=2,R_{\text{reach target}}=2,w_{\text{nav}}=1$
@@
$R_{\text{manip}}=$ $\delta(w_{\text{manip shaping}}R_{\text{manip shaping}})$
@@
where $\delta$ is defined as:
$\delta=\begin{cases}1,&\text{if the end effector haven't reached the target region}\\ 0,&\text{otherwise}\end{cases}$
$\gamma$ is defined as:
$\gamma=\begin{cases}1,&\text{if the robot finish the task at the current step}\\ 0,&\text{otherwise}\end{cases}$
$d_{\text{closet}}$ is the closest distance the end-effector reached before towards the specified region of the target object, $d_{\text{current}}$ is the current distance between the end-effector and the specified region of the target object, $P$ is the task progress. For all tasks, we use $w_{\text{manip shaping}}=0.02,R_{\text{finish task}}=20$. For opening door/fridge tasks, we use $w_{\text{progress}}=80$. For cleaning table task, we use $w_{\text{progress}=100}$. For opening door (pull) and opening fridge task, we provide additional $w_{grasp}=2$ for grasping the door knob or the edge of the fridge.
@@
where $d_{\text{ee moved}}$ is the distance end-effector moved at the current step, and $\gamma$ is defined as:
$\gamma=\begin{cases}1,&\text{if current action failed }\\ 0,&\text{otherwise}\end{cases}$
Current action fails if the robot collide into semenatic objects in the apartment in simulation, such as wall or table. For all tasks, we use $R_{\text{step penalty}}=-0.01,w_{\text{ee moved}}=-0.01,R_{\text{invalid action}}=-0.01$

### v3

@@ V-A Training and Evaluation Details
V-B Performance of HarmonicMM in Simulation
@@
Our experiment results demonstrate that HarmonicMM achieves an average success rate of 47% across all tasks (As in Table I). Specifically, our model attains a high success rate of 80% for the task of Opening Door (Push) in unseen houses, and achieves average progress rates of 76%, 70%, and 46% for Cleaning Table, Opening Door (Pull), and Opening Fridge, respectively, indicating significant progress in each task. This achievement is especially impressive given the complex, long-horizon nature of these tasks.
@@
The success rates across tasks vary due to how the robot interacts with its environment. Opening Door (Push) has a high success rate as it’s less affected by environmental factors: the robot can use the space created by the opened door to continue the task. However, Opening Door (Pull) and Opening Fridge are more challenging. These tasks reduce the available space as they progress, especially in narrow areas, leading to lower success rates despite significant progress. Cleaning Table involves complex navigation due to obstacles and furniture around the table, requiring intricate movement patterns. This complexity accounts for its lower performance compared to door-related tasks. For Opening Fridge, the task’s difficulty arises from the smaller size of fridge doors and the need for precise end-effector placement. The often confined placement of fridges in rooms adds to the challenge, affecting the robot’s ability to navigate and manipulate effectively.
Figure 5: Cleaning Table in Simulation
@@
Comparison with the previous approaches. We hypothesize that jointly learning and executing navigation and manipulation for mobile manipulation tasks achieves better performance than considering them disjointly. To validate this, we compared our method with two common approaches widely used by the community. For a fair comparison, we use the same network architecture stated in Sec IV and the same hyperparameters for baselines.
@@
One common practice, which we note as Two-Stage baseline, in embodied AI for mobile manipulation ([17, 35, 50]) is to decompose the tasks into separate stages of manipulation and navigation. In this work, our Two-Stage baseline uses a roughly similar high-level structure as OVMM[17] in which the robot navigates to the target object and, after reaching the target, switches to the manipulation. While this approach works for a variety of pick-and-place tasks, it fails for more complex tasks, such as those presented in our benchmark. Table I shows the very low success rate and progress of the Two-Stage baseline.
@@
Another common approach is executing manipulation and navigation alternatively. At every step controller invokes separately obtained manipulation or navigation skills. We adapted ReLMoGen [11] to support the embodiment of the Stretch robot. This baseline outputs the target end-effector pose, target base pose in the current robot frame, and choice over navigation or manipulation at every step. Although this approach achieves much higher results than the Two-Stage baseline, there is still a significant 17.6% (37% relative) drop in success rate compared to HarmonicMM.
HarmonicMM is Efficient. Confirming our hypothesis, HarmonicMM not only boosts performance but also enhances efficiency. In table cleaning, our method produces efficient motion of extending arms while approaching the table, as shown in Fig. 5. Table I shows that HarmonicMM makes 32.2% more progress towards completing the task at each step compared to the baselines on Cleaning Table, 113.4% on Opening Door (Push), and 27.6% on Opening Door (Pull).
V-C HarmonicMM transfers to the real world.
TABLE III: Real World Evaluation
@@
Cleaning Table $66.6\%$ 12
Opening Door (Pull) $60.0\%$ 15
Opening Door (Push) $70.0\%$ 10
TABLE IV: Ablation Study on Opening Door (Pull)
Opening Door (Pull) Success Rate Progress Reach Door
No Pretrained DINOv2 $0\%$ $0\%$ $66.3\%$
Nav Cam Only $27.3\%$ $47.9\%$ $93\%$
Manip Cam Only $0\%$ $0.235\%$ $95\%$
[…]
Our learned policies were evaluated in a real apartment for Opening Door (Pull), Opening Door (Push), and Cleaning Table tasks, showing promising outcomes as in Table III and Fig.  4. The Opening Fridge task is not evaluated since the magnetic seal of the fridge is too strong for our robot to pull. To compensate for the simulation’s simplified magnetic gripper grasping, a heuristic function for doorknob grasping was implemented in the real world, effectively bridging the simulation-to-reality gap. When our controller initiates a grasp action, the heuristic function is invoked to grasp the doorknob.
@@
Our controller successfully pulled the door fully open in 9 out of 15 attempts in three different rooms. Interestingly, we observed that HarmonicMM exhibited two distinct styles of pulling the door to adapt to the layout differences in each room. When there was enough open space around the door, it continued pulling until the door was fully opened, as shown in the 1st row in Fig. 4. However, when the door was located next to a wall, the agent first pulled the door to start the opening process and then switched to pushing to fully open it (2nd row, Fig. 4). These emergent behaviors showcase the spatial reasoning and scene-understanding abilities of our controller. Similarly, our controller successfully pushed doors fully open 7 times in 4 different rooms out of 10 trials.
@@
For the cleaning table task, we placed the table in four different locations across two rooms, with evenly distributed paper pieces on it for the robot to clean. Our robot navigated to the table and continuously moved its end-effector to clean the entire table, as shown in Fig.  4. Out of 12 trials, our controller successfully cleaned $66.6\%$ of the table’s surface area in eight instances.
@@
Figure 6: Improvement over Range: The improvement HarmonicMM obtained over ReLMoGen grows as the distance increases.
V-D Ablation Study of HarmonicMM
@@
More efficient in longer horizon tasks. we analyzed the success rate of both HarmonicMM and ReLMoGen[11] in the Cleaning Table and Opening Door (Push) tasks across varying initial distances from the target object of interest (aka. table and door), ranging from $1\sim 4$. As in Fig.  6, we observed that the improvement of HarmonicMM over the baseline becomes larger as the initial distance to the target object increases. This trend underscores HarmonicMM’s enhanced efficiency, particularly in tasks with longer horizons, where the coordinated integration of navigation and manipulation is pivotal in successful task completion.
Proprioception. Our comparison of HarmonicMM against a proprioception-less variant in door-opening tasks illustrates the proprioception significance (in Table II). We also observe $-33.7\%$ / $-20\%$ success rate drop, and $-8.3\%$ / $-26.2\%$ progress drop for Cleaning Table and Opening Fridge without proprioception. The absence of proprioceptive feedback notably diminishes performance, especially in Opening Door (Pull), where precise end-effector positioning is necessary for pre-grasping the knob. This effect is less pronounced but still present in Opening Door (Push), emphasizing proprioception’s overall contribution to task performance and efficiency, even when precise manipulation is not as critical.
Multi-View Visual Observation. We had initial experiments with a single-camera setup showing a significant performance drop compared with multi-camera ones (provided in Table IV). This indicates that complex tasks require simultaneous visual input for both navigation and manipulation, validating our approach of integrated navigation and manipulation.
Pretrained Visual Encoder We substituted the DINOv2 encoder with a trainable CNN. This modification led to a significant drop in success rate and progress (Tab.IV), underscoring the critical role of a pretrained visual encoder
Transformer Variant. We adapted Skill Transformer[51] and evaluated it on the most complicated task Open Door (Pull) (As in Table  IV). It performs poorly, likely due to the low-sample efficiency of training a large transformer with a ResNet18 encoder from scratch (following the original paper), which is unsuitable for our online RL setting. The original paper trains different low-level skills separately and uses a skill transformer for coordinating skills, which is significantly different from our task setting.
VI Conclusion
@@
In this work, we developed a simulation for complex mobile manipulation tasks, ranging from opening doors/fridges to cleaning tables in daily scenes, using ProcThor[12]. Our HarmonicMM enables robots to solve these tasks using only RGB visual observation and proprioception of the robot through end-to-end learning. Our pipeline significantly outperforms previous baselines in simulation and has successfully transferred to real-world apartments with novel layouts, without any fine-tuning. Code for our benchmark and our method will be released.
Limitation: Despite its success, our HarmonicMM faces limitations related to the kinematic and physical capabilities of the robot, restricting its ability to handle more dynamic real-world tasks, such as lifting heavy objects.
@@
Acknowledgements. We would like to thank the Thor Team: Winson Han, Eli VanderBilt, and Alvaro Herrasti, for their support in setting up the simulations. Special thanks to Winson Han for his expertise and assistance in the visualization of this work.
Appendix
VII Visual Augmentation
Our applied augmentations include: ColorJitter: Adjusts brightness (0.4), contrast (0.4), saturation (0.2), and hue (0.05); GaussianBlur: Applies a blur effect with a kernel size of (5,9) and sigma range of (0.1,2); RandomResizedCrop: Resizes the input with a scale range of (0.9,1); RandomPosterize: Reduces color depth, applied with varying bits (7,6,5,4) and a probability of 0.2 for each setting; RandomAdjustSharpness: Enhances or reduces the sharpness with a factor of 2, applied with a probability of 0.5.
VIII Network Architecture
Our RGB observations initially get processed by DinoV2 model, generating $768\times 16\times 16$ features which pooled to a reduced dimension of $768\times 7\times 7$. The processed features are passed through separate visual encoders. This results in two distinct sets of $16\times 7\times 7$ latent visual features corresponding to each camera view. The key hyperparameters are as follows:
@@ VIII Network Architecture
RGB Input $384\times 224$
Pretrained Visual Encoder dinov2_vits14
$E_{Nav}$ 2 Conv(1x1, 16 channels)
$E_{Manip}$ 2 Conv(1x1, 16 channels)
@@
Policy($\pi$) 1 FC with 512 units
Value($v$) 1 FC with 512 units
@@
latent feature dim $7*7*16*2+5=1573$
IX RL training Hyperparameters
@@
X Reward Function
In this section, we provide the formulation of our reward and hyperparameters for different tasks in Sec IV-D.
@@ X Reward Function
$R_{\text{nav}}=\delta w_{\text{nav shaping}}max(d_{\text{closest}}-d_{\text{current}},0)+\gamma R_{\text{reach target}}$
The variable $\delta$ is $1$ if the robot hasn’t reached the target position and 00 otherwise. Similarly, $\gamma$ is 1 if the robot reaches the target at the current step, and 0 otherwise. $d_{\text{closest}}$ is the closest distance robot reached before, $d_{\text{current}}$ is the current distance between robot and target object. For all tasks, we use $w_{\text{nav shaping}}=2,R_{\text{reach target}}=2,w_{\text{nav}}=1$
@@
$R_{\text{manip}}=$ $\delta w_{\text{manip shaping}}R_{\text{manip shaping}}$
@@
The variable $\delta$ is $1$ if the end effector has not reached the target region, and 00 otherwise. Similarly, $\gamma$ is 1 if the robot completes the task at the current step, and 0 otherwise.
$d_{\text{closest}}$ is the closest distance the end-effector reached before towards the specified region of the target object, $d_{\text{current}}$ is the current distance between the end-effector and the specified region of the target object, $P$ is the task progress. For all tasks, we use $w_{\text{manip}}=1,w_{\text{manip shaping}}=0.02,R_{\text{finish task}}=20$. For opening door/fridge tasks, we use $w_{\text{progress}}=80$. For the cleaning table task, we use $w_{\text{progress}=100}$. For opening the door (pull) and opening the fridge task, we provide an additional $w_{grasp}=2$ for grasping the door knob or the edge of the fridge. We consider the grasp successful if the agent has issued the grasp action and the distance between the end-effector and object is smaller than $0.2$.
@@
where $d_{\text{ee moved}}$ is the distance end-effector moved at the current step, and $\gamma$ is 1 if the current action failed, and 0 otherwise. Current action fails if the robot collides with semantic objects in the apartment in simulation, such as a wall or table. For all tasks, we use $R_{\text{step penalty}}=-0.01,w_{\text{ee moved}}=-0.01,R_{\text{invalid action}}=-0.01,w_{\text{efficiency}}=1$
Hardware Platform: Our simulation and real-world experiment are conducted with Stretch Robot [43] and we use RealSense D455 for our RGB observation.
Future Works: In future work, we aim to enhance the capabilities of HarmonicMM by integrating more complex and dynamic tasks into our task suite. We aim to explore the potential of HarmonicMM in environments with even longer task horizons and more varied challenges.